OpenStudy (anonymous):

Find the area of the region bounded the curves y=e^2 (x>0), y=e^(-2x) (x<0), and x-axis, x=-1 to x=1.

OpenStudy (anonymous):

isnt the second integral suppose to be e^x ? p.s thanks

OpenStudy (unklerhaukus):

what were the equations for the curves ?

OpenStudy (anonymous):

srry: y=e^x (x>0), y=e^(-2x) (x<0), and x-axis, x=-1 to x=1

OpenStudy (anonymous):

Can someone confirm my ans. I got 4.9128.

OpenStudy (unklerhaukus):